# supplementary

### v1

@@ IV Experiments
We validate RWM through a comprehensive set of experiments across diverse robotic systems, environments, network architectures. The experiments are designed to assess the accuracy and robustness of RWM, evaluate its architectural and training design choices, and demonstrate its effectiveness across diverse robotic tasks and in real-world deployment combined with MBPO-PPO. We start the analysis by looking into the autoregressive prediction accuracy and robustness of the world model on ANYmal D learned with simulation data induced by a velocity tracking policy. The observation and action spaces of the world model are detailed in Table S1 and Table S3. We then compare various network architectures and the error induced across diverse robotic environments and tasks to demonstrate the generality of RWM. And finally, we learn a policy in RWM with the proposed MBPO-PPO and demonstrate the applicability and robustness of the method on an ANYmal D hardware.
IV-A Autoregressive Trajectory Prediction
@@
The capability of a world model to maintain high fidelity during autoregressive rollouts is critical for effective planning and policy optimization. To evaluate this aspect, we analyze the autoregressive prediction performance of RWM using trajectories collected from ANYmal D hardware. The control frequency of the robot is at $50\,Hz$. The model is trained with history horizon $M=32$ and forecast horizon $N=8$. Further details on the network architecture and training parameters are summarized in Sec. A-B1 and Sec. A-C1, respectively. The autoregressive trajectory predictions by RWM are visualized in Fig. 3.
@@
Figure 3: Autoregressive trajectory prediction by RWM. Solid lines represent ground truth trajectories, while dashed lines denote predicted state evolution. The control frequency of the robot is at $50\,Hz$. The model is trained with history horizon $M=32$ and forecast horizon $N=8$. Predictions commence at $t=32$ using historical observations, with future observations predicted autoregressively by feeding prior predictions back into the model.
At time $t=32$, the world model initiates predictions based on historical observations. From this point onward, future observations are predicted autoregressively using its prior outputs, simulating the deployment scenario in downstream applications. In Fig. 3, the predicted observations are compared against the ground truth trajectories.
@@
The results demonstrate that RWM exhibits a remarkable alignment between predicted and ground truth trajectories across all observed variables. This consistency persists over extended rollouts, showcasing the model’s ability to mitigate compounding errors—a critical challenge in long-horizon predictions. Despite the accumulated shifts at the far end of the prediction, the variables maintain close fidelity to the actual trajectories, with minimal deviations, even in stochastic and dynamically complex environments.
The proposed GRU-based autoregressive mechanism demonstrates superior predictive accuracy. The historical context encoded in the hidden states allows RWM to effectively capture partially observable dynamics, a capability evident in the accurate long-term predictions. This performance is further attributed to the dual-autoregressive mechanism introduced in Sec. III-B, which stabilizes predictions despite the short forecast horizon employed during training. This robust predictive capability has profound implications for downstream tasks. A comparison of state evolution between the RWM prediction and the ground truth simulation is illustrated in Fig. LABEL:fig:prediction_visualization (bottom). Both environments are initialized in the same state and propagated using the same policy to ensure a fair evaluation of the world model’s predictive accuracy.
The visualization highlights the ability of RWM to maintain consistency in trajectory predictions over long horizons, even beyond the training forecast horizon. This robustness is pivotal for stable policy learning and deployment, as discussed further in Sec. IV-E. The results underscore the capability of RWM to produce accurate long-horizon predictions, validating its applicability to real-world robotic tasks.
IV-B Robustness under Noise
@@
A critical challenge in training world models is their ability to generalize under noisy conditions, particularly when predictions rely on autoregressive rollouts. Even small deviations from the training distribution can cascade into untrained regions, causing the model to hallucinate future trajectories. These issues are magnified in autoregressive settings due to the compounding nature of prediction errors. To assess the robustness of RWM, we analyze its performance under Gaussian noise perturbations applied to both observations and actions. We compare the results with an MLP-based baseline also trained autoregressively with the same history and forecast horizon, as shown in Fig. 4, where yellow curves denote the relative prediction error $e$ for RWM, and grey curves represent the MLP baseline.
Figure 4: Autoregressive trajectory prediction error under Gaussian noise. Yellow curves denote RWM at varying noise levels, demonstrating consistent robustness and lower error accumulation across forecast steps. Grey curves represent the MLP baseline, which exhibits significantly higher error accumulation and reduced robustness to noise.
@@
The results indicate a clear advantage of RWM over the MLP baseline across all noise levels. As forecast steps increase, the relative prediction error of the MLP model grows significantly, diverging more rapidly than RWM. In contrast, RWM demonstrates superior stability, maintaining lower prediction errors even under high noise levels. This robustness can be attributed to the dual-autoregressive mechanism introduced in Sec. III-B, which ensures stability in long-horizon predictions. The inner autoregression updates the GRU hidden states during the historical context window, while the outer autoregression feeds back predictions during the forecast horizon. This design minimizes the accumulation of errors by continually refining the state representation toward long-term predictions, even in the presence of noisy inputs.
IV-C Dual-autoregressive Mechanism
The choice of history horizon $M$ and forecast horizon $N$ plays a critical role in the training and performance of RWM. The history horizon $M$ governs the amount of historical information the model utilizes to infer the current state and predict future observations. The forecast horizon $N$ defines the extent of supervised forward prediction steps during training. These two hyperparameters directly impact the autoregressive prediction accuracy and training time, as shown in Fig. 5.
Figure 5: Ablation study on the history horizon $M$ and forecast horizon $N$ in RWM. The heatmap on the left shows the relative autoregressive prediction error, with darker colors indicating higher errors. Models trained with larger history horizons $M$ exhibit lower errors, although the improvements plateau beyond a certain point. Forecast horizon $N$ has a significant impact, with longer horizons leading to better long-term prediction accuracy due to exposure to extended rollouts during training. The heatmap on the right illustrates training time, with darker colors representing longer durations. Increasing $N$ significantly raises training time due to sequential computation, while shorter horizons (e.g., $N=1$, teacher-forcing) enable faster training but result in poor prediction accuracy.
The heatmap on the left in Fig. 5 shows the relative autoregressive prediction error $e$ under different combinations of $M$ and $N$. Models trained with a longer history horizon $M$ consistently exhibit lower prediction errors, demonstrating the importance of providing sufficient historical context to capture the underlying dynamics. However, the influence of $M$ plateaus beyond a certain point, indicating diminishing returns for very large history horizons. Forecast horizon $N$, on the other hand, plays a decisive role in improving long-term prediction accuracy. Increasing $N$ during training leads to better performance in autoregressive rollouts, as it encourages the model to learn representations robust to compounding errors over extended prediction horizons. This improvement comes at the cost of increased training time, as shown in the heatmap on the right. Larger $N$ values require sequential computation during training due to the autoregressive nature of the process, significantly lengthening the training duration.
IV-D Generality across Robotic Environments
@@
To assess the generality and robustness of RWM across a diverse range of robotic environments, we compare its performance with several baseline methods, including MLP, recurrent state-space model (RSSM) [18, 33, 14, 34], and transformer-based architectures [48, 50]. These baselines represent widely adopted approaches in dynamics modeling and policy optimization. All models are given the same context during training and evaluation. Their training parameters are detailed in Sec. A-B2. The relative autoregressive prediction errors $e$ for these models across various environments are shown in Fig. 6. The tasks span manipulation scenarios as well as quadruped and humanoid locomotion tasks, allowing for a comprehensive evaluation of the models. In addition, we highlight the importance of the autoregressive training introduced in Sec. III-B by including both RWM trained with teacher-forcing (RWM-TF) and autoregressive training (RWM-AR), demonstrating the significant performance gains achieved by the latter.
@@
The results highlight the superiority of RWM trained with autoregressive training (RWM-AR), which consistently achieves the lowest prediction errors across all environments. The performance gap between RWM-AR and the baselines is especially pronounced in complex and dynamic tasks, such as velocity tracking for legged robots, where accurate long-horizon predictions are critical for effective control. The comparison also reveals that RWM-AR significantly outperforms its teacher-forcing counterpart (RWM-TF), underscoring the importance of autoregressive training in mitigating compounding prediction errors over long rollouts. We additionally visualize the imagination rolled out by RWM-AR compared with the ground truth simulation in Fig. LABEL:fig:prediction_visualization and Fig. 7.
Figure 7: Autoregressive imagination of RWM and ground-truth simulation across diverse robotic systems. For each environment, the top row showcases the RWM autoregressively predicting future trajectories in imagination. The second row visualizes the ground truth evolution in simulation. The visualized coordinate and arrow markers denote the predicted and measured end-effector pose and base velocity, respectively.
One of the key factors driving this performance is RWM’s dual-autoregressive mechanism, which ensures robustness in long-horizon predictions by sequentially incorporating historical context and feeding back its own predictions during rollouts. In contrast, models such as RSSM and transformers often exhibit error accumulation due to their reliance on shorter training horizons or weaker temporal representations. This capability is particularly evident in tasks with dynamically complex trajectories, such as legged locomotion, where slight inaccuracies in prediction can cascade into large errors over time. Importantly, the autoregressive training approach further amplifies the robustness of RWM by simulating the conditions encountered during deployment. While teacher-forcing training (RWM-TF) performs adequately in environments with shorter prediction horizons, it struggles with long rollouts where compounding errors dominate. Autoregressive training (RWM-AR), on the other hand, exposes the model to its own predictions during training, enabling it to better handle error propagation and produce reliable trajectories over extended horizons.
@@
Note that the baselines are trained using teacher forcing as they are traditionally implemented. However, the proposed autoregressive training framework is architecture-agnostic and can also be applied to baseline models. When trained with autoregressive training, RSSM achieves a performance comparable to the proposed GRU-based architecture. Nevertheless, we opt for the GRU-based model due to its simplicity and computational efficiency. On the other hand, training transformer architectures with autoregressive training does not scale effectively, as the multi-step gradient propagation in autoregressive forecasting leads to GPU memory constraints, limiting their practicality for this approach. These results demonstrate that RWM, when combined with autoregressive training, achieves robust and generalizable performance across diverse robotic tasks. Its ability to consistently outperform baseline architectures highlights the effectiveness of its design in addressing the challenges of long-horizon prediction and error accumulation.
IV-E Policy Learning and Hardware Transfer
Using MBPO-PPO, we train a goal-conditioned velocity tracking policy for ANYmal D leveraging RWM. At the beginning of each imagination rollout, a target velocity is sampled alongside the initial observation retrieved from the replay buffer ${\mathcal{D}}$. The policy selects an action based on the sampled observation and velocity command. RWM then predicts the next observation based on the observation-action pair, and the reward is computed from this transition. This process continues recursively over the rollout horizon, with the collected imagination trajectories being used to update the policy. After each iteration, the updated policy interacts with the real environment, collecting additional data for the replay buffer, which is subsequently used to refine RWM.
The policy’s observation and action spaces are detailed in Sec. A-A1, and its architecture is described in Sec. A-B3. Reward formulations are provided in Sec. A-A2, while training parameters are summarized in Sec. A-C2.
Note that learning policies with long imagination rollouts from a learned world model is especially challenging due to the inherent compounding errors in the autoregressive nature of the predictions arising from two key factors: The world model’s reliance on previous predictions to infer future observations, which introduces drift over time. And the policy’s dependence on predicted observations to compute actions, amplifying errors in the generated trajectories. We compare MBPO-PPO with two baselines: Short-Horizon Actor-Critic (SHAC) [45] and DreamerV3 [34]. SHAC employs a first-order gradient-based method that propagates gradients through the world model to optimize the policy. Dreamer, in contrast, integrates a latent-space dynamics model with an actor-critic framework, emphasizing sample efficiency and robustness in continuous control tasks. The comparative results of model error and policy performance are shown in Fig. 8.
@@ IV-E Policy Learning and Hardware Transfer
Figure 8: Model error and policy mean reward for the ANYmal D velocity tracking task with MBPO-PPO. The policy is trained using estimated rewards computed from predicted observations by RWM. Ground truth rewards, visualized with solid lines, are reported by the simulator for evaluation purposes only.
@@
Figure 8 (top) illustrates the model error $e$ during policy optimization. While MBPO-PPO demonstrates a significant reduction in model error over training, stabilizing at a consistently low value, SHAC struggles with high and fluctuating model error throughout the process. Its reliance on first-order gradients for optimization is not well-suited for discontinuous dynamics, such as those encountered in legged locomotion, where system behavior changes drastically due to varying contact patterns. The resulting inaccurate gradients lead to suboptimal policy updates, producing chaotic robot behaviors during training. These chaotic behaviors, in turn, generate low-quality training data for updating RWM, exacerbating model inaccuracies. This feedback loop causes a compounding failure mode where both the policy and the model collapse. Dreamer exhibits lower model error than SHAC but still significantly higher error compared to MBPO-PPO. Although Dreamer effectively leverages its latent-space dynamics model for policy optimization, its reliance on shorter planning horizons during training limits its ability to handle long-horizon dependencies, particularly in stochastic environments. As a result, Dreamer encounters moderate compounding errors during policy learning, which hinder its convergence to optimal behaviors. While Dreamer’s model error stabilizes faster than SHAC, it remains insufficiently accurate for effective sim-to-real transfer.
On the bottom plot illustrating rewards $r$, the predicted rewards (dashed lines) for MBPO-PPO slightly overshoot the ground truth rewards (solid lines) during the early and middle stages of training. This overshooting occurs because the policy optimization process amplifies any small inaccuracies in the world model’s predicted rewards, particularly during the policy’s initial improvement phases. The world model, although accurate, learns a slightly optimistic approximation of the reward function, which the policy leverages during optimization. This leads to an overestimation of expected rewards in the learned policy. Despite this overshooting, the predicted rewards converge closer to the ground truth as training progresses, indicating that the model-predicted rewards remain sufficiently accurate to guide effective policy learning. Importantly, MBPO-PPO ultimately achieves a high ground truth reward, demonstrating that the overshooting does not hinder its convergence or real-world applicability.
In contrast, SHAC fails to converge entirely. The policy learned under SHAC generates erratic and ineffective robot behaviors, which not only prevent meaningful policy improvement but also undermine the accuracy of the learned world model. This is particularly detrimental for challenging dynamics like legged locomotion, where accurate policy optimization is crucial for producing stable and effective behaviors. Dreamer demonstrates partial convergence, achieving higher rewards compared to SHAC but significantly lagging behind MBPO-PPO. The policies learned using Dreamer exhibit smoother and more stable behaviors than SHAC but fail to match the robustness and performance of the policies optimized with RWM. This discrepancy highlights Dreamer’s limited capacity to generalize to the intricate and discontinuous dynamics of legged locomotion tasks.
@@
To evaluate the robustness of the learned policy, we deploy it on ANYmal D hardware in a zero-shot transfer setup. SHAC and Dreamer fail to produce a deployable policy due to its collapse during training. However, as shown in Fig. LABEL:fig:prediction_visualization, the policy learned using MBPO-PPO demonstrates reliable and robust performance in tracking goal-conditioned velocity commands and maintaining stability under external disturbances, such as unexpected impacts and terrain conditions. The success of MBPO-PPO in hardware deployment is a direct result of the high-quality trajectory predictions generated by RWM, which enable accurate and effective policy optimization. Videos showcasing the robustness of the policy in hardware, including its responses to external disturbances, are available in our supplementary materials. These results underline the effectiveness of RWM and MBPO-PPO in enabling robust and scalable policy deployment for real-world robotic systems.
V Limitations

### v4

@@ 4 Experiments
We validate RWM through a comprehensive set of experiments across diverse robotic systems, environments, and network architectures. The experiments are designed to assess the accuracy and robustness of RWM, evaluate its architectural and training design choices, and demonstrate its effectiveness across diverse robotic tasks in Isaac Lab [mittal2023orbit] and in real-world deployment combined with MBPO-PPO. We start the analysis by looking into the autoregressive prediction accuracy and robustness of the world model learned with simulation data induced by a velocity tracking policy. The observation and action spaces of the world model are detailed in Table S2 and Table S4. We then compare various network architectures and the error induced across diverse robotic environments and tasks to demonstrate the generality of RWM. And finally, we learn a policy in RWM with the proposed MBPO-PPO and demonstrate the applicability and robustness of the method on ANYmal D [hutter2016anymal] and Unitree G1 hardware.
4.1 Autoregressive Trajectory Prediction
@@
The capability of a world model to maintain high fidelity during autoregressive rollouts is critical for effective planning and policy optimization. To evaluate this aspect, we analyze the autoregressive prediction performance of RWM using trajectories collected from ANYmal D hardware. The control frequency of the robot is at 50Hz50\,Hz. The model is trained with history horizon $M=32$ and forecast horizon $N=8$. Further details on the network architecture and training parameters are summarized in Sec. A.2.1 and Sec. A.3.1, respectively. The autoregressive trajectory predictions by RWM are visualized in Fig. 3(a).
(a) Autoregressive trajectory prediction by RWM.
(b) Prediction error under Gaussian noise.
@@
Figure 3: (Left) Solid lines represent ground truth trajectories, while dashed lines denote predicted state evolution. Predictions commence at $t=32$ using historical observations, with future observations predicted autoregressively by feeding prior predictions back into the model. (Right) Yellow curves denote RWM at varying noise levels, demonstrating consistent robustness and lower error accumulation across forecast steps. Grey curves represent the MLP baseline, which exhibits significantly higher error accumulation and reduced robustness to noise.
@@
The results demonstrate that RWM exhibits a remarkable alignment between predicted and ground truth trajectories across all observed variables. This consistency persists over extended rollouts, showcasing the model’s ability to mitigate compounding errors—a critical challenge in long-horizon predictions. This performance is attributed to the dual-autoregressive mechanism introduced in Sec. 3.2, which stabilizes predictions despite the short forecast horizon employed during training. A comparison of state evolution between the RWM prediction and the ground truth simulation is illustrated in Fig. 1 (bottom). The visualization highlights the ability of RWM to maintain consistency in trajectory predictions over long horizons, even beyond the training forecast horizon. This robustness is pivotal for stable policy learning and deployment, as discussed further in Sec. 4.4.
It is notable that the choice of history horizon $M$ and forecast horizon $N$ plays a critical role in the training and performance of RWM. Our ablation study in Sec. A.4.1 reveals that, while extending both $M$ and $N$ improves accuracy, practical considerations of computational cost necessitate careful tuning of these hyperparameters to achieve optimal performance.
4.2 Robustness under Noise
@@
A critical challenge in training world models is their ability to generalize under noisy conditions, particularly when predictions rely on autoregressive rollouts. Even small deviations from the training distribution can cascade into untrained regions, causing the model to hallucinate future trajectories. To assess the robustness of RWM, we analyze its performance under Gaussian noise perturbations applied to both observations and actions. We compare the results with an MLP-based baseline also trained autoregressively with the same history and forecast horizon, as shown in Fig. 3(b), where yellow curves denote the relative prediction error $e$ for RWM, and grey curves represent the MLP baseline.
@@
The results indicate a clear advantage of RWM over the MLP baseline across all noise levels. As forecast steps increase, the relative prediction error of the MLP model grows significantly, diverging more rapidly than RWM. In contrast, RWM demonstrates superior stability, maintaining lower prediction errors even under high noise levels. This robustness can be attributed to the dual-autoregressive mechanism introduced in Sec. 3.2, which ensures stability in long-horizon predictions. This design minimizes the accumulation of errors by continually refining the state representation toward long-term predictions, even in the presence of noisy inputs.
4.3 Generality across Robotic Environments
@@
To assess the generality and robustness of RWM across a diverse range of robotic environments, we compare its performance with several baseline methods, including MLP, recurrent state-space model (RSSM) [hafner2019learning, hafner2019dream, hafner2020mastering, hafner2023mastering], and transformer-based architectures [chen2021decision, reed2022generalist]. These baselines represent widely adopted approaches in dynamics modeling and policy optimization. All models are given the same context during training and evaluation. Their training parameters are detailed in Sec. A.2.2. The relative autoregressive prediction errors $e$ for these models are shown in Fig. 4. The tasks span manipulation scenarios as well as quadruped and humanoid locomotion tasks, allowing for a comprehensive evaluation of the models. In addition, we highlight the importance of the autoregressive training introduced in Sec. 3.2 by including both RWM trained with teacher-forcing (RWM-TF) and autoregressive training (RWM-AR), demonstrating the significant performance gains achieved by the latter.
@@
The results highlight the superiority of RWM trained with autoregressive training (RWM-AR), which consistently achieves the lowest prediction errors across all environments. The performance gap between RWM-AR and the baselines is especially pronounced in complex and dynamic tasks, such as velocity tracking for legged robots, where accurate long-horizon predictions are critical for effective control. The comparison also reveals that RWM-AR significantly outperforms its teacher-forcing counterpart (RWM-TF), underscoring the importance of autoregressive training in mitigating compounding prediction errors over long rollouts. We additionally visualize the imagination rolled out by RWM-AR compared with the ground truth simulation in Fig. 1 and Fig. S9.
@@
Note that the baselines are trained using teacher forcing as they are traditionally implemented. However, the proposed autoregressive training framework is architecture-agnostic and can also be applied to baseline models. When trained with autoregressive training, RSSM achieves a performance comparable to the proposed GRU-based architecture. Nevertheless, we opt for the GRU-based model due to its simplicity and computational efficiency. On the other hand, training transformer architectures with autoregressive training does not scale effectively, as the multi-step gradient propagation in autoregressive forecasting leads to GPU memory constraints, limiting their practicality for this approach. These results demonstrate that RWM, when combined with autoregressive training, achieves robust and generalizable performance across diverse robotic tasks.
4.4 Policy Learning and Hardware Transfer
Using MBPO-PPO, we train a goal-conditioned velocity tracking policy for ANYmal D and Unitree G1 leveraging RWM. The policy’s observation and action spaces are detailed in Sec. A.1.1, and its architecture is described in Sec. A.2.3. Reward formulations are provided in Sec. A.1.2, while training parameters are summarized in Sec. A.3.2. We compare MBPO-PPO with two baselines: Short-Horizon Actor-Critic (SHAC) [xu2022accelerated] and DreamerV3 [hafner2023mastering]. SHAC employs a first-order gradient-based method that propagates gradients through the world model to optimize the policy. Dreamer integrates a latent-space dynamics model with an actor-critic framework, emphasizing sample efficiency and robustness in continuous control tasks.
@@ 4.4 Policy Learning and Hardware Transfer
Figure 5: Model error and policy mean reward for the ANYmal D (left) and Unitree G1 (right) velocity tracking task with MBPO-PPO. The policy is trained using estimated rewards computed from predicted observations by RWM. Ground truth rewards, visualized with solid lines, are reported by the simulator for evaluation purposes only.
@@
Figure 5 illustrates the model error $e$ during policy optimization. While MBPO-PPO demonstrates a significant reduction in model error over training, SHAC struggles with high and fluctuating model error throughout the process. Its reliance on first-order gradients for optimization is not well-suited for discontinuous dynamics, such as those encountered in legged locomotion, where system behavior changes drastically due to varying contact patterns. The resulting inaccurate gradients lead to suboptimal policy updates, producing chaotic robot behaviors during training. These chaotic behaviors, in turn, generate low-quality training data for updating RWM, exacerbating model inaccuracies. Although Dreamer effectively leverages its latent-space dynamics model for policy optimization, its reliance on shorter planning horizons during training limits its ability to handle long-horizon dependencies, particularly in stochastic environments. As a result, Dreamer encounters moderate compounding errors during policy learning, which hinder its convergence to optimal behaviors.
On the right plot of rewards $r$, predicted rewards (dashed) from MBPO-PPO initially overshoot the ground truth (solid) due to the policy exploiting small inaccuracies in the model’s optimistic estimates. As training progresses, predictions align more closely with ground truth, remaining accurate enough to guide effective learning. In contrast, SHAC fails to converge, producing unstable behaviors that degrade both policy and model quality. Dreamer demonstrates partial convergence, achieving higher rewards compared to SHAC but significantly lagging behind MBPO-PPO.
@@
To evaluate the robustness of the learned policies, we deploy them on ANYmal D and Unitree G1 hardware in a zero-shot transfer setup. SHAC and Dreamer fail to produce a deployable policy due to its collapse during training. However, as shown in Fig. 1, the policy learned using MBPO-PPO demonstrates reliable and robust performance in tracking goal-conditioned velocity commands and maintaining stability under external disturbances, such as unexpected impacts and terrain conditions. The success of MBPO-PPO in hardware deployment is a direct result of the high-quality trajectory predictions generated by RWM, which enable accurate and effective policy optimization. Videos showcasing the robustness of the policies on hardware, including their responses to external disturbances, are available on our webpage. These results underline the effectiveness of RWM and MBPO-PPO in enabling robust and scalable policy deployment for real-world robotic systems.
5 Limitations